# Scalable Multi-Agent Path Finding using Collision-Aware Dynamic Alert Mask and a Hybrid Execution Strategy

### v1

@@ Abstract
Multi-agent pathfinding (MAPF) remains a critical problem in robotics and autonomous systems, where agents must navigate shared spaces efficiently while avoiding conflicts. Traditional centralized algorithms that have global information, such as Conflict-Based Search (CBS), provide high-quality solutions but become computationally expensive in large-scale scenarios due to the combinatorial explosion of conflicts that need resolution. Conversely, distributed approaches that have local information, particularly learning-based methods, offer better scalability by operating with relaxed information availability, yet often at the cost of solution quality. To address these limitations, we propose a hybrid framework that combines decentralized path planning with a lightweight centralized coordinator. Our framework leverages reinforcement learning (RL) for decentralized planning, enabling agents to adapt their planning based on minimal, targeted alerts–such as static conflict-cell flags or brief conflict tracks–that are dynamically shared information from the central coordinator for effective conflict resolution. We empirically study the effect of the information available to an agent on its planning performance. Our approach reduces the inter-agent information sharing compared to fully centralized and distributed methods, while still consistently finding feasible, collision-free solutions–even in large-scale scenarios having higher agent counts.
@@ 1 Introduction
Multi-Agent Path Finding (MAPF) addresses the fundamental challenge of computing collision-free trajectories for multiple agents navigating a shared environment. Its effective solving is crucial for deploying a wide array of multi-agent systems, from automated warehouses and robotic swarms to autonomous vehicle coordination (Sharon et al. 2015). Despite its practical significance, MAPF is computationally demanding, classified as NP-hard in its general form, which can render traditional centralized search methods intractable as the number of agents or the complexity of the environment increases (Ren et al. 2025; Sartoretti et al. 2019).
@@
Extensive research in MAPF has explored various coordination paradigms, each with different implications for information availability and system performance. Centralized approaches, such as Conflict-Based Search (CBS) (Sharon et al. 2015) and its efficiency-focused variant ICBS (Boyarski et al. 2015), typically assume full observability of the environment and agent states, helping them generate optimal plans. However, as the number of agents or the size of the environment increases, joint planning and conflict resolution become computationally expensive. Additionally, requiring agents to share full information raises privacy concerns, limiting applicability in confidential settings motivating strategies that reduce reliance on global information. Early decoupled methods like $M^{*}$ (Wagner and Choset 2011) begin with individual plans and only coordinate agents in conflict, offering scalability but often sacrificing optimality or completeness. More recently, Multi-Agent Reinforcement Learning (MARL) has emerged to address coordination and partial observability in dynamic settings. For example, PRIMAL (Sartoretti et al. 2019) trains agents to plan using partial views, learning implicit coordination, while methods like FOLLOWER (Skrynnik et al. 2024) reduce explicit communication by relying on global heuristic maps. These decentralized methods improve scalability but may reduce solution quality and leave conflicts unresolved due to limited information.
@@
To address these limitations, we propose a novel hybrid MAPF framework that combines decentralized planning with a lightweight centralized coordinator, focusing on minimizing inter-agent information sharing while maintaining solution feasibility. Our findings show that minimal, targeted alerts are sufficient, reducing the total information load by an estimated $\sim$93% compared to a continuous-observation distributed paradigm (see Appendix A.1). In our approach, agents primarily rely on decentralized, reinforcement learning-based neural network planners that operate on local observations (e.g., position coordinates), eliminating the need for global information as well as local egocentric maps. A central coordinator oversees agent trajectories, intervening selectively by dynamically sharing targeted information to prompt localized re-planning when conflicts are anticipated. This work’s contributions are centered on this selective coordination strategy. We introduce (1) a novel hybrid framework with this on-demand alert mechanism and provide (2) a comprehensive empirical evaluation demonstrating that our planning policy generalizes robustly from a simple training regimen to larger, more complex environments. Our evaluation is guided by two key research questions: (RQ1) Given the observation constraints of a decentralized setup, can an effective MAPF algorithm be created with one agent knowing nothing about other agents? If not, what information must it need at a minimum? and (RQ2) How does the proposed hybrid method compare to leading alternative search- and learning-based approaches in terms of performance, solution quality, and scalability?
@@ 2.1 Multi‑Agent Path Finding
Let $G=(V,E)$ be an undirected graph, where $V$ is the set of vertices (grid cells) and $E\subseteq V\times V$ is the set of edges connecting adjacent cells. A team of $n$ agents $A=\{a_{1},\dots,a_{n}\}$ must move from start vertices $s_{i}\in V$ to goal vertices $g_{i}\in V$, where $(s_{i},g_{i})\neq(s_{j},g_{j}),\forall i\neq j:i,j\in\{1,...,n\}$. Time is discretized into steps $t=0,1,2,\dots$, and at each step, an agent may either move along an edge or wait. A path for agent $a_{i}$ is a sequence $\pi_{i}=(v^{i}_{0},v^{i}_{1},\dots,v^{i}_{T_{i}})$ with $v^{i}_{0}=s_{i}$ and $v^{i}_{T_{i}}=g_{i}$. A solution to MAPF is $\Pi=\{\pi_{1},\dots,\pi_{n}\}$, and it is collision‑free if for all $i\neq j$ and all $t$, $v^{i}_{t}\neq v^{j}_{t}$ (vertex collision free) and $(v^{i}_{t},v^{i}_{t+1})\neq(v^{j}_{t+1},v^{j}_{t})$ (edge collision free). The primary goal in standard MAPF is typically to find a path set $\Pi$ that is collision-free (Sharon et al. 2015). Common efficiency objectives include minimizing the makespan, $\max_{i}T_{i}$, minimizing the sum of individual completion times (sum-of-costs), $\sum_{i}T_{i}$, or minimizing the number of collisions.
An important consideration in defining a MAPF problem instance arises from agents potentially reaching their targets at different time steps, thus requiring a clear definition of how an agent behaves after it has reached its target $g_{i}$ at time $T_{i}$, but while other agents may still be en route. Two common settings are used to address this scenario (Stern et al. 2019a). The first setting, typically referred to as stay at target, considers that an agent, upon reaching its goal, remains at that vertex until all other agents in the team have also reached their respective targets. This waiting agent continues to occupy its target vertex, potentially causing conflicts if another agent’s plan requires traversing or occupying that vertex at any time $t\geq T_{i}$ up to the completion time of the entire solution. The second setting, disappear at target, assumes an agent is effectively removed from the environment immediately upon reaching its target. Consequently, its path concludes at time $T_{i}$, and it poses no further collision risk nor occupies any vertex for $t>T_{i}$. In our work, we adopt the second setting, disappear at target. This setting is particularly relevant for applications like drone delivery, where an agent’s task is considered complete upon arrival at a target location.
@@ 2.2 Coordination Paradigms
As formalized by Sharon et al. (Sharon et al. 2015), solution approaches to MAPF problems can be categorized based on their coordination strategy. We extend this by focusing on three key elements: state observability (global vs. local), communication (allowed vs. none), and control (centralized vs. decentralized). In centralized paradigms, a global planner with full observability of the entire state controls all agents, and communication is implicit through this central controller. By contrast, other approaches grant agents local control over their actions. Within this category, we draw a key distinction based on communication: In distributed approaches, agents that plan their own paths are allowed to explicitly exchange information, such as local observations, intended paths, or goals, with other agents to achieve cooperative behavior. In decentralized MAPF, as defined in this work, inter-agent communication during execution is eliminated. Agents plan and act entirely independently, relying only on their own local information. Finally, hybrid frameworks, like the one we propose, combine these elements. They typically employ a central coordination mechanism that has full observability to detect conflicts and selectively intervene, while agents otherwise plan in a decentralized manner. In our work, we adopt a hybrid framework with a customized information sharing mechanism as described in Section 3.
@@ 2.3 Literature Review
Search-based Methods: Research on MAPF has evolved from centralized to more scalable methods. Early advancements include $M^{*}$, which dynamically couples agent searches upon conflict for optimality and completeness (Wagner and Choset 2011), and Increasing Cost Tree Search (ICTS), which employs cost allocation and then checks for conflict-free solutions from sets of single-agent paths, yielding significant speedups over joint-space search (Sharon et al. 2013). A pivotal development was CBS, where agents independently plan paths using $A^{*}$, and a centralized detector resolves conflicts by branching with new constraints for involved agents, ensuring optimality (Sharon et al. 2015). To enhance scalability while preserving optimality, variants such as ICBS integrate cardinality heuristics and meta-agent merging to accelerate convergence (Boyarski et al. 2015). For very large instances, suboptimal repair strategies like Large Neighborhood Search (LNS) generate initial solutions (often via prioritized planning) and iteratively repair colliding agent subsets under a centralized control loop (Li et al. 2022). However, these methods face scalability challenges as the number of agents increases and often depend on extensive global information for coordination.
Learning-based Methods: In parallel, learning‐based approaches address partial observability and communication constraints. PRIMAL and PRIMAL2 train decentralized policies via a combination of imitation learning from an expert centralized planner and reinforcement learning (PRIMAL2 enhancing local observations for better coordination), effectively implementing a centralized training and decentralized execution (CTDE) paradigm (Sartoretti et al. 2019; Damani et al. 2021b). SCRIMP introduces a transformer‐based communication module enabling agents with limited fields of view to coordinate effectively, achieving decentralized coordination without a central execution-time coordinator (Wang et al. 2023). To function in real‐world settings, each agent must carry onboard sensing, e.g., RGB‐D cameras or LiDAR scanners, and implement an inter-agent communication protocol for goal disclosure, which introduces additional computational overhead and raises privacy concerns. While offering adaptability and decentralized execution, learning-based methods can require significant training data, may struggle with generalization, and often provide weaker guarantees on solution quality or completeness when relying purely on local information.
Hybrid Methods: Hybrid planning‐learning paradigms seek the strengths of both worlds. Authors in (Skrynnik et al. 2024) introduce FOLLOWER, which involves each agent using a congestion‐aware $A^{*}$ planner to determine its subgoals, then executing a fully decentralized RL policy for lifelong replanning. More recently, hybrid frameworks like LNS2+RL, introduced by Li et al., have emerged, augmenting LNS by employing MARL for early, localized conflict resolution, then relying on efficient search-based planning for final refinement (Wang et al. 2025). These existing hybrid approaches, while innovative, can involve complex integrations and may still face trade-offs between optimality, scalability, or specific information dependencies (like global congestion maps (Skrynnik et al. 2024)).
@@
To enable a structured comparison of diverse MAPF techniques (search-based, learning-based, and hybrid), we adopt the four‑stage MAPF pipeline: S1 (Agent Planning), S2 (Collision Detection), S3 (Collision Avoidance Policy), and S4 (Agent Replanning). Within this structure, CBS uses decentralized A* search for S1/S4 and centralized high level solver for collision resolution in S2/S3 for optimality (Sharon et al. 2015). LNS employs a centralized loop for all four stages, managing path generation, detection, repair, and subset replanning, thereby trading optimality for scalability (Li et al. 2022). Learning-based planners like PRIMAL and SCRIMP typically use local observations for decentralized S1-S4 execution, coordinating without a central controller (Sartoretti et al. 2019; Wang et al. 2023); however, practical deployments can require sensors (eg., camera, LiDAR etc.) and communication, adding computational overhead and increase sensing cost. Hybrid methods such as FOLLOWER use global congestion maps with $A^{*}$ search for S1, then decentralized RL policy for S2-S4, often without further communication (Skrynnik et al. 2024). LNS2+RL applies centralized LNS for S1-S3, while S4 uses a hybrid policy (MARL or prioritized planning) for subsets, balancing key trade-offs (Wang et al. 2025). Refer to Table 2 in the Appendix Section A.2 for a detailed categorization of these methodologies.
@@
Together, these works highlight key limitations of existing MAPF paradigms: centralized methods face scalability and privacy concerns; purely learning-based approaches compromise solution quality; and current hybrid solutions often still depend on significant centralized communication or coordination overhead with higher sensing cost. The hybrid framework introduced in Section 3, which is the focus of our work, is specifically designed to mitigate these limitations. It strategically reduces inter-agent information sharing to maintain solution feasibility, thereby addressing the aforementioned information management challenges while also aiming to preclude the additional sensing costs and computational overhead common in learning-based methods.
@@ 3 Methodology
3.1 Approach Summary
We propose a hybrid coordination framework that leverages decentralized RL-based path planning together with centralized collision detection and control. While conceptually similar to the high-level loop of Conflict-Based Search (CBS), our framework differs significantly in its resolution step: it does not build a constraint tree or perform backtracking, instead issuing targeted alerts that prompt heuristic-based replanning from the local RL agent. The proposed method’s four-stage MAPF pipeline is detailed below.

### v2

# Towards Information-Optimized Multi-Agent Path Finding: A Hybrid Framework with Reduced Inter-Agent Information Sharing

@@ Abstract
Multi-agent pathfinding (MAPF) remains a critical problem in robotics and autonomous systems, where agents must navigate shared spaces efficiently while avoiding conflicts. Traditional centralized algorithms with global information provide high-quality solutions but scale poorly in large-scale scenarios due to the combinatorial explosion of conflicts. Conversely, distributed approaches that have local information, particularly learning-based methods, offer better scalability by operating with relaxed information availability, yet often at the cost of solution quality. In realistic deployments, information is a constrained resource: broadcasting full agent states and goals can raise privacy concerns, strain limited bandwidth, and require extra sensing and communication hardware, increasing cost and energy use. We focus on the core question of how MAPF can be solved with minimal inter-agent information sharing while preserving solution feasibility. To this end, we present an information-centric formulation of the MAPF problem and introduce a hybrid framework, IO-MAPF, that integrates decentralized path planning with a lightweight centralized coordinator. In this framework, agents use reinforcement learning (RL) to plan independently, while the central coordinator provides minimal, targeted signals, such as static conflict-cell indicators or short conflict trajectories, that are dynamically shared to support efficient conflict resolution. We introduce an Information Units (IU) metric to quantify information use and show that our alert-driven design achieves $2\times$ to $23\times$ reduction in information sharing, compared to the state-of-the-art algorithms, while maintaining high success rates, demonstrating that reliable MAPF is achievable under strongly information-restricted, privacy-preserving conditions. We demonstrate the effectiveness of our algorithm using both simulation and hardware experiments.
@@ 1 Introduction
Multi-Agent Path Finding (MAPF) addresses the fundamental problem of computing collision-free trajectories for multiple agents navigating a shared environment. Solving it effectively is crucial for a wide range of real-world applications, ranging from the deployment of automated robotic swarms in warehouses to performing autonomous vehicle coordination. Despite its practical significance, solving MAPF is computationally demanding, classified as NP-hard in its general form, which can render traditional centralized search methods intractable as the number of agents or the complexity of the environment increases (Ren et al. 2025; Sartoretti and others 2019).
@@
The MAPF literature has explored various coordination paradigms, each with different implications for information availability and system performance. Centralized approaches, such as Conflict-Based Search (CBS) (Sharon et al. 2015) and its efficiency-focused variant ICBS (Boyarski et al. 2015), typically assume full knowledge of the environment and agent states and compute collision-free paths based on this global information available. However, as the number of agents or the size of the environment increases, joint planning and conflict resolution become computationally expensive. Additionally, requiring agents to share full information raises privacy concerns, limiting applicability in confidential settings and imposing additional compute requirements in real-world deployments, motivating strategies that reduce reliance on global information. Early decoupled methods like $M^{*}$ (Wagner and Choset 2011) begin with individual plans and only coordinate agents in conflict, offering scalability but often sacrificing optimality or completeness. More recently, Multi-Agent Reinforcement Learning (MARL) has emerged to address coordination and partial observability in dynamic settings. For example, PRIMAL (Sartoretti and others 2019) trains agents to plan using partial views, learning implicit coordination, while methods like FOLLOWER (Skrynnik et al. 2024) reduce explicit communication by relying on global heuristic maps. These decentralized methods improve scalability but may reduce solution quality and leave conflicts unresolved due to limited information.
@@
To address these limitations, we first formulate the MAPF problem from an information-centric perspective and with the aim of reducing information exchange under this setting, we propose Information-Centric MAPF (IC–MAPF), a novel hybrid MAPF framework that combines decentralized planning with a lightweight centralized coordinator, that minimizes inter-agent information sharing while maintaining solution feasibility. Our findings show that minimal, targeted alerts are sufficient, reducing the total information load by an estimated $2\times$ to $23\times$ compared to the state-of-the-art communication efficient algorithms (see Table 3). In our approach, agents primarily rely on decentralized, reinforcement learning-based neural network planners that operate on local observations (e.g., position coordinates), eliminating the need for global information as well as local egocentric maps. A central coordinator oversees agent trajectories, intervening selectively by dynamically sharing targeted information to prompt localized re-planning when conflicts are anticipated. This work’s contributions are centered on this selective coordination strategy. We introduce (1) a information-centric MAPF formulation and identify the information utilized by various MAPF strategies, (2) IC–MAPF - a novel hybrid framework with an on-demand alert mechanism that substantially reduces the amount of global information an agent must access for successful planning, and (3) a new metric, Information Units (IU), to quantify the information available to each agent during planning.
Our evaluation is guided by two key research questions: (RQ1) Given the observation constraints of a decentralized setup, can an effective MAPF algorithm be created with one agent knowing nothing about other agents? If not, what information must it need at a minimum? and (RQ2) How does the proposed hybrid method compare to leading alternative search- and learning-based approaches in terms of performance, solution quality, and scalability?
@@ 2.1 Multi‑Agent Path Finding
Let $G=(V,E)$ be an undirected graph, where $V$ is the set of vertices (grid cells) and $E\subseteq V\times V$ is the set of edges connecting adjacent cells. A team of $n$ agents $A=\{a_{1},\dots,a_{n}\}$ must move from start vertices $s_{i}\in V$ to goal vertices $g_{i}\in V$, where $(s_{i},g_{i})\neq(s_{j},g_{j}),\forall i\neq j:i,j\in\{1,...,n\}$. Time is discretized into steps $t=0,1,2,\dots$, and at each step, an agent may either move along an edge or wait. A path for agent $a_{i}$ is a sequence $\pi_{i}=(v^{i}_{0},v^{i}_{1},\dots,v^{i}_{T_{i}})$ with $v^{i}_{0}=s_{i}$ and $v^{i}_{T_{i}}=g_{i}$. A solution to MAPF is $\Pi=\{\pi_{1},\dots,\pi_{n}\}$, and it is collision‑free if for all $i\neq j$ and all $t$, $v^{i}_{t}\neq v^{j}_{t}$ (vertex collision free) and $(v^{i}_{t},v^{i}_{t+1})\neq(v^{j}_{t+1},v^{j}_{t})$ (edge collision free). The primary goal in standard MAPF is typically to find a path set $\Pi$ that is collision-free (Sharon et al. 2015). Common efficiency objectives include minimizing the makespan, $\max_{i}T_{i}$, minimizing the sum of individual completion times (sum-of-costs), $\sum_{i}T_{i}$.
An important modeling choice in MAPF concerns how agents are treated after reaching their goals at potentially different times. Two common settings are used (Stern et al. 2019a). In the stay-at-target setting, an agent remains at its goal vertex after arrival and continues to occupy that location until all agents have finished, potentially creating conflicts for others. In contrast, the disappear-at-target setting removes an agent from the environment immediately upon goal arrival, eliminating any further interactions. In this work, we adopt the stay-at-target setting, which is appropriate for persistent agents operating in shared environments.
@@ 2.2 Coordination Paradigms
As formalized by Sharon et al. (Sharon et al. 2015), solution approaches to MAPF problems can be categorized based on their coordination strategy. We extend this by focusing on three key elements: state observability (global vs. local), communication (allowed vs. none), and control (centralized vs. decentralized). In centralized paradigms, a global planner with full observability of the entire state controls all agents, and communication is implicit through this central controller. By contrast, other approaches grant agents local control over their actions. Within this category, we draw a key distinction based on communication: In distributed approaches, agents that plan their own paths are allowed to explicitly exchange information, such as local observations, intended paths, or goals, with other agents to achieve cooperative behavior. In decentralized MAPF, as defined in this work, inter-agent communication during execution is eliminated. Agents plan and act entirely independently, relying only on their own path information. Finally, hybrid frameworks, like the one we propose, combine these elements. They typically employ a central coordination mechanism that has full observability to detect conflicts and selectively intervene, while agents otherwise plan in a decentralized manner. In our work, we adopt a hybrid framework with a customized information sharing mechanism as described in Section 3.
@@ 2.3 Literature Review
Search-based Methods: Classical MAPF research progressed from centralized optimal solvers to more scalable search frameworks. Early methods such as $M^{*}$ dynamically couple agents only when conflicts arise (Wagner and Choset 2011), while ICTS allocates costs to agents and searches for conflict-free combinations of single-agent paths, reducing joint-space complexity (Sharon et al. 2013). A major milestone was CBS, which plans independently for each agent using $A^{*}$ and resolves conflicts by branching on constraints, guaranteeing optimality (Sharon et al. 2015). Its variants, including ICBS, further accelerate search through cardinality heuristics and meta-agent merging (Boyarski et al. 2015). For large-scale problems, suboptimal approaches like LNS generate an initial solution (often via prioritized planning) and iteratively repair conflicting subsets (Li et al. 2022). PIBT (Okumura et al. 2022) introduces adaptive prioritization for local agent movements, and LaCAM (Okumura 2023) combines low-level constraint extraction with high-level location-sequence search. Despite their efficiency gains, these methods still face scalability limits with growing agent counts and rely heavily on global information.
Learning-based Methods: Learning-based MAPF methods focus on handling partial observability and limited communication through decentralized policies trained under the centralized training and decentralized execution (CTDE) paradigm. PRIMAL and PRIMAL2 combine imitation learning from an expert centralized solver with reinforcement learning (PRIMAL2 further improving local observations), enabling scalable decentralized execution (Sartoretti and others 2019; Damani et al. 2021). SCRIMP employs a transformer-based communication module that supports coordination among agents with restricted field-of-view (FOV), allowing decentralized conflict avoidance without a central coordinator (Wang et al. 2023). SYLPH (He et al. 2025) enables agents to infer their Social Value Orientation (SVO), a measure of selfishness or altruism, based on situational context, and adapt their behavior toward the most influential agent in the system. ALPHA (He et al. 2024) blends accurate proximal sensing with fuzzy global cues to guide agents beyond purely local reasoning, mitigating local myopia. SIGMA (Liao et al. 2025) uses sheaf-theoretic local consensus to capture geometric cross-dependencies among agents. In real deployments, each agent must carry onboard sensors (e.g., RGB-D or LiDAR) and maintain communication channels for goal disclosure, introducing extra computation and potential privacy risks. While adaptable and fully decentralized, learning-based methods often require extensive training, may generalize poorly, and provide weaker guarantees on solution quality or completeness when relying solely on local information.
Hybrid Methods: Hybrid planning–learning approaches aim to combine the strengths of both paradigms. FOLLOWER (Skrynnik et al. 2024) uses congestion-aware $A^{*}$ planning to generate subgoals, followed by decentralized RL for replanning. DCC (Ma et al. 2021b) and EPH (Tang et al. 2024) augment local FOV sensing with selective inter-agent communication, while DHC (Ma et al. 2021a) exchanges information with $K$ nearest neighbors to improve coordination. EPH further employs ensemble-based inference and switches to fast $A^{*}$ when no agents appear in FOV. More recent systems such as LNS2+RL (Wang et al. 2025) combine LNS with MARL for early conflict resolution before refining solutions via search. Although effective, these hybrid designs often require complex integrations and specific information dependencies, and typically trade optimality for scalability.
@@
To enable a structured comparison across search-based, learning-based, and hybrid MAPF methods, we adopt a four-stage pipeline: S1 (Agent Planning), S2 (Collision Detection), S3 (Collision Avoidance Policy), and S4 (Agent Replanning). Within this view, CBS performs decentralized $A^{*}$ for S1/S4 and uses a centralized high-level solver for S2/S3, ensuring optimality. LNS centralizes all four stages, path generation, conflict detection, repair, and subset replanning, trading optimality for scalability. Learning-based approaches rely on local observations to decentralize S1–S4, though real deployments typically require onboard sensing (e.g., cameras, LiDAR) and inter-agent communication, increasing cost and computation. Hybrid methods occupy the middle ground: FOLLOWER uses global congestion maps with $A^{*}$ for S1 and decentralized RL for S2–S4, while LNS2+RL combines centralized LNS for S1–S3 with a hybrid RL/prioritized-planning strategy for S4. A detailed categorization is provided in Table 6 (Appendix Section A.4).
@@
Overall, existing MAPF paradigms face key trade-offs: centralized methods struggle with scalability and privacy; purely learning-based methods can weaken solution quality; and current hybrids still rely on substantial centralized coordination or costly sensing. The hybrid framework, IC–MAPF, introduced in Section 3 is designed to overcome these limitations by reducing inter-agent information sharing while preserving feasibility, thereby avoiding the communication and sensing overhead typical of learning-based systems.
@@ 3 Methodology
3.1 Information-Centric MAPF
Algorithm Category Information Used by Agent $i$
CBS/ EECBS/ LaCAM/ LaCAM*/ LNS/ LNS2 Centralized (Search-Based) $\mathcal{I}_{i}^{\text{central}}(t)=(\mathcal{M},\,\mathcal{S}_{i}=(s_{i}^{t},g_{i}^{t},\pi_{i|t_{0}:t_{\max}}),\,\mathcal{O}_{j}=(s_{j}^{t},g_{j}^{t},\pi_{j|t_{0}:t_{\max}})_{j\neq i})$
MRCDRL / AMPPO-PP Centralized (Learning-Based) $\mathcal{I}_{i}^{\text{central}}(t)=(\mathcal{M},\,\mathcal{S}_{i}=(s_{i}^{t},g_{i}^{t},\pi_{i|t_{0}:t_{\max}}),\,\mathcal{O}_{j}=(s_{j}^{t},g_{j}^{t},\pi_{j|t_{0}:t_{\max}})_{j\neq i})$
PIBT / CS-PIBT / Distributed (Search-Based) $\mathcal{I}_{i}^{\text{distributed}}(t)=(\mathcal{M}_{i}^{k},\,\mathcal{S}_{i}=(s_{i}^{t},g_{i}^{t},\pi_{i|t_{s}:t_{r}}),\,\mathcal{O}_{j}=\{(s_{j}^{t},g_{j}^{t},\pi_{j|t_{s}:t_{r}})\}_{j\in\mathcal{N}_{i}(t)})$
SACHA / PRIMAL / SCRIMP Distributed (Learning-Based) $\mathcal{I}_{i}^{\text{distributed}}(t)=(\mathcal{M}_{i}^{k},\,\mathcal{S}_{i}=(s_{i}^{t},g_{i}^{t},\pi_{i|t_{s}:t_{r}}),\,\mathcal{O}_{j}=\{(s_{j}^{t},g_{j}^{t},\pi_{j|t_{s}:t_{r}})\}_{j\in\mathcal{N}_{i}(t)})$
FOLLOWER Decentralized (Learning-Based) $\mathcal{I}_{i}^{\text{decentral}}(t)=(\mathcal{M}_{i}^{k},\,\mathcal{S}_{i}=(s_{i}^{t},g_{i}^{t},\pi_{i|t_{s}:t_{r}}))$
MA-CBS/ M*/ HiPP-MAPP/ DCC/ DHC/ EPH/ IC–MAPF Hybrid $\mathcal{I}_{i}^{\text{hybrid}}(t)=(\mathcal{M}_{i},\,\mathcal{S}_{i}=(s_{i}^{t},g_{i}^{t},\pi_{i|t_{s}:t_{r}}),\,\mathcal{O}_{j}=\{(s_{j}^{t},g_{j}^{t},\pi_{j|t_{s}:t_{r}})\}_{j\in\mathcal{N}_{i}(t)},\mathcal{E}_{ij})$
Table 1: Categorization of MAPF algorithms by information availability, based on the unified model $\mathcal{I}_{i}(t)=(\mathcal{M}_{i},\mathcal{S}_{i},\mathcal{O}_{j})$.
In an MAPF problem, each agent $i\in\{1,\dots,N\}$ relies on a set of information components to safely navigate from its start to its goal. This includes environmental information, represented by the entire map or a subset of the map $\mathcal{M}_{i}=(V,E,O)$ of vertices, edges, and static obstacles; self information, consisting of the agent’s current state $s_{i}^{t}\in V$, goal $g_{i}^{t}\in V$, and its augmented policy over a time interval $\{t_{s},...,t_{r}\}$ given by $\pi_{i|t_{s}:t_{r}}$, which can be compactly denoted by $\mathcal{S}_{i}=(s_{i}^{t},g_{i}^{t},\pi_{i|t_{s}:t_{r}})$; and some amount of other-agent information, specifically their current positions, goal positions and the immediate intended actions following their individual policies, $\mathcal{O}_{j}=(s_{j}^{t},g_{j}^{t},\pi_{j|t_{s}:t_{r}})_{j\in\mathcal{N}_{i}(t)}$, where $\mathcal{N}_{i}(t)$ denotes the set of agents in the neighborhood of agent $i$, required to avoid vertex and edge collisions. Collectively, the information accessible to agent $i$ at any timestep t, is written as $\mathcal{I}_{i}(t)=\bigl(\mathcal{M}_{i},\,\mathcal{S}_{i},\mathcal{O}_{j}\bigr).$ While this set captures the general information required for MAPF, different algorithmic paradigms make use of $\mathcal{I}_{i}(t)$ in distinct ways. In centralized algorithms, a global solver gathers the full map information and the full joint information $\{s_{i}^{t},\,g_{i}^{t}\}_{i=1}^{N}$ and computes conflict-free paths $\{\pi_{i}\}_{i=1}^{N}$; during execution, each agent requires only $\mathcal{I}_{i}^{\text{central}}(t)=(\mathcal{M}_{i},\mathcal{S}_{i},\mathcal{O}_{j})$ where $\mathcal{M}_{i}=\mathcal{M}$, and $t_{s}=t_{0}$ and $t_{r}=t_{max}$, where $\{t_{0},...,t_{max}\}$ denotes the entire time-interval required for the agent to traverse from its start to goal positions. Note that the central coordinator under this setting requires the information $O_{j},j\neq i$ of all the other agents, including the neighboring agents.
In contrast, distributed algorithms require each agent to make decisions based solely on locally available information, leading to $\mathcal{I}_{i}^{\text{distributed}}(t)=\bigl(\mathcal{M}_{i}^{k},\mathcal{S}_{i},\mathcal{O}_{j}\bigr),$ where $\mathcal{M}_{i}$ may be a $k\times k$ subgrid of the map, often referred to as the FOV, $\mathcal{S}_{i}$ is the agent’s self information and $\mathcal{O}_{j}$, with $j\in\mathcal{N}_{i}$ is the neighboring agents’ information which is either sensed or exchanged.
On the other hand, in decentralized algorithms, each agent only has access to the environment information and self-information, using which they independently plan out paths. So the information in these strategies can be denoted as $\mathcal{I}_{i}^{\text{decentral}}(t)=\bigl(\mathcal{M}_{i}^{k},\mathcal{S}_{i}\bigr).$
Finally, hybrid algorithms combine global and local reasoning: a central module resolves large-scale couplings or computes partial plans, while agents use local information for real-time coordination, yielding $\mathcal{I}_{i}^{\text{hybrid}}(t)=\bigl(\mathcal{M}_{i},\,\mathcal{S}_{i},\mathcal{O}_{j},\mathcal{E}_{ij}\bigr),$ where additional information $\mathcal{O}_{j}$ related to other agents is shared to an agent only under the occurrence of specific events $\mathcal{E}_{ij}$, like collisions or deadlocks. Thus, although all MAPF settings require the same foundational information, they differ fundamentally in how much of it is shared, how widely it is distributed, and which components are used to compute safe and coordinated multi-agent behavior. To illustrate this, the information exchange across various categories of well known MAPF algorithms is presented in Table 1.
3.2 Proposed Approach for IC–MAPF
To address this problem of information-centric MAPF, we treat collision resolution as a mechanism for controlling when additional information about other agents needs to be shared. At a high level, a hybrid four-stage strategy can be built. In S1 (Decentralized Path Planning), each agent independently computes a plan from its start to its goal using information about the map and static obstacles. In S2 (Centralized Collision Detection), a coordinating module examines the collection of agent plans over time and identifies conflicts (vertex or edge conflicts). In S3 (Collision Control), a central control policy interprets each detected conflict and specifies a set of constraints to resolve the conflict. Finally, in S4 (Replanning), one of the agents involved in the conflict replans its path following the constraints from S3.
By varying what constraints are set for replanning, we obtain a family of tiered replanning strategies (S4.1–S4.4) that trade off information usage against solution completeness. (S4.1) yield-based local coordination, where a single agent performs a short detour to a nearby parking location, waits, and then rejoins its original plan using only immediate (typically 1 to 3 hop distance) local occupancy information; (S4.2) static replanning, where a small segment of one agent’s plan is recomputed while treating a fixed set of conflict cells as forbidden and keeping all other plans unchanged; (S4.3) dynamic replanning, where a small segment is replanned while treating short prefixes of other agents’ plans as time-varying obstacles over a finite horizon; and (S4.4) local joint planning, where a tightly coupled subset of agents is replanned jointly over a bounded window while all remaining agents are treated as fixed reservations. Together, these tiers provide a controlled way to gradually increase the amount of information used for coordination only when simpler, lower inter-agent information-based repairs fail. We provide completeness and soundness properties of these individual strategies and the aggregated strategy we employed in this work in Appendix A.2. The proposed methodology of the four-stage MAPF pipeline is detailed below.